# The ShareLM Collection and Plugin: Contributing Human-Model Chats for the Benefit of the Community

### v1

@@ 1 Introduction
Despite being a cornerstone for LLM development and research, mechanisms for openly collecting and sharing human conversations and feedback are still underdeveloped. In the meantime, models developed by for-profit companies collect user-model conversations via their APIs to be used to further train their own models (Ouyang et al., 2022), leaving the open-source and research community far behind. The development process of these “closed models” is not always transparent, and so are their data and data collection pipelines. These all make developing platforms and tools for collecting human-model conversations a high priority.
@@ 3.2 System Architecture
The plugin works by identifying certain elements in the web page XML, according to the chat platform in use. Currently, the plugin supports Gradio [^3] https://www.gradio.app/, a web interface for various demos including chats, and ChatUI [^4] https://huggingface.co/docs/chat-ui/index, a web interface for chats. Those two were chosen due to their popularity, e.g., they are in frequent use in Huggingface Spaces [^5] https://huggingface.co/spaces and the ChatBot Arena (see §7). Nevertheless, adding support to new web platforms is easy [^6] We were informed by private correspondence that external contributors are planning to extend the support to other interfaces.
@@ 4.1 Terms of Use
To activate the plugin after installation, the user needs to confirm the terms of use. The terms of use are available through the plugin popup (see §4.3), or the recording banner while in a supported demo (see next §4.2). We ask the users to avoid sharing conversations with identifying/sensitive content (names, e-mail addresses, etc.), as the content of the conversations will be publicly released. The full terms are available in the plugin repository.
@@ 4.2 The Recording Banner
The main role of the recording banner is to inform the user their conversations are recorded. In addition, it can be used to pause the conversation sharing. Clicking on the "Click Here to Stop Sharing" button will turn off the conversations collection. This is useful when conducting a conversation with identifying information that should be kept private.
@@ 4.3 The Popup
Under the conversations table, we include a frequently asked questions section, to answer common questions regarding the plugin. There, we address questions regarding privacy (e.g., Will it be possible to identify me by my conversations?), license (Would you share the dataset? With what license?), ownership (How can I ask to remove all my conversations from the dataset?) among others.
@@ 8 Conclusions and Future Work
We would also like to extend the plugin for more commonly used chat interfaces, possibly for closed models too.
@@
Currently, we are not supporting the collection of conversations from ’closed’ models, as we wanted to concentrate on open source models. However, as many people still prefer the closed models for their day-to-day use, we miss a lot of data. We will consider adding support for closed models in the future.

### v2

@@ 1 Introduction
Despite being a cornerstone for LLM development and research, mechanisms for openly collecting and sharing human conversations and feedback are still underdeveloped. In the meantime, models developed by for-profit companies collect user-model conversations via their APIs to be used to further train their own models (Ouyang et al., 2022), leaving the open-source and research community far behind. The development process of these “closed models” is not always transparent, and so are their data and data collection pipelines. These all make developing platforms and tools for collecting human-model conversations a high priority (Don-Yehiya et al., 2024).
@@ 3.2 System Architecture
The plugin works by identifying certain elements in the web page XML, according to the chat platform in use. Currently, the plugin supports Gradio [^3] https://www.gradio.app/, a web interface for various demos including chats, ChatUI [^4] https://huggingface.co/docs/chat-ui/index, a web interface for chats, and ChatGPT. Those were chosen due to their popularity, e.g., Gradio and ChatUI are in frequent use in Huggingface Spaces [^5] https://huggingface.co/spaces and the ChatBot Arena (see §7). Nevertheless, adding support to new web platforms is easy [^6] We were informed by private correspondence that external contributors are planning to extend the support to other interfaces.
@@
Figure 3: The conversation collection is paused. Clicking on the "Click here to start sharing" button will start the conversation’s recording.
@@ 4.1 Terms of Use
To activate the plugin after installation, the user needs to confirm the terms of use. The terms of use are available through the plugin popup (see §4.3), or the recording banner while in a supported demo (see next §4.2). We ask the users to avoid sharing conversations with identifying/sensitive content (names, e-mail addresses, etc.), as the content of the conversations will be publicly released. The full terms are available in the plugin repository and in App. §A.
@@ 4.2 The Recording Banner
The main role of the recording banner is to inform the user their conversations are recorded. In addition, it can be used to pause the conversation sharing. Clicking on the "Click Here to Stop Sharing" button will turn off the conversations collection (see Fig. 3). This is useful when conducting a conversation with identifying information that should be kept private.
@@ 4.3 The Popup
Under the conversations table, we include a frequently asked questions section, to answer common questions regarding the plugin (see Fig. 5). There, we address questions regarding privacy (e.g., Will it be possible to identify me by my conversations?), license (Would you share the dataset? With what license?), ownership (How can I ask to remove all my conversations from the dataset?) among others.
@@
Figure 5: The frequently asked questions section (in the popup window). Provides answers to common questions regarding the plugin.
@@ Acknowledgments
Appendix A Terms of Use
Inspired by the release of the ChatGPT, the open-source community recently began to develop open access models with increased transparency about their development. The next challenge for democratizing large language models is data.
This extension collects the conversations you are having with open large language models (“chat-bots”). By using this extension, you are giving your permission to contribute your conversations’ content (both your side of the conversation, and the model’s) for creating an open-license chat-bot conversations dataset, a valuable resource for the open-source community. The conversations will be released with the most permissive license that is allowed by the specific model. This dataset will be a valuable resource for both model developers and researchers. Specifically, we plan to use this dataset to study and improve the nature of human- model interaction.
The extension supports a couple of chat-bots demos, mostly within Huggingface Spaces (https://huggingface.co/spaces). You will see a banner on the top of the demo page indicating it. You can choose not to share a particular conversation by clicking the ‘do not share’ button. As an additional precaution, the conversations are not posted to the database immediately. You can see the conversations from the last 24 hours in the extension popup window and remove them. To stop sharing your conversations permanently, please disable or remove the extension. Note that removing the extension does not delete the conversations you have already made.
Along with the conversation’s content, we are collecting the URL (to identify the model), GMT time and an anonymous user-id. Optionally, you can fill some demographic data (age, location, gender) and rate your satisfaction. We are not collecting any identifying metadata (such as IP address, local time, browser type, etc.). However, it is possible that you will be identified by the content of your conversations. Therefore, please avoid sharing conversations with Identifying/sensitive content (names, e-mail addresses, etc.), as the content of your conversations will be publicly released. If you accidentally shared the content of a conversation you prefer to keep private, please fill the contact form so we will remove it (available in the extension popup). You can ask to remove all your conversations at any time, but please note that after the dataset was already released it is very likely that someone has already downloaded and saved an old version of it. You are encouraged to use this form also for reporting conversations that are copyrighted, defamatory, threatening to others, violating of others’ privacy, or that you view as harmful if released.
Please be advised that this extension is independently developed by us, and while we have put our best efforts into ensuring a smooth experience, it’s important to note that there might be bugs or unforeseen issues. Your feedback is valuable to us, so please feel free to report any issues you may encounter.
The research is conducted by Shachar Don-Yehiya, Leshem Choshen and Omri Abend at the Hebrew University.
For more questions, please contact us at shareLM.project@gmail.com.
Participation is from age 18 and over only.
Participation is voluntary. Thank you for your contribution!